Uniform Motion
Dave and Jane drove a total of 935 miles in 15.2 hours. Dave drove the first part of the trip and averaged 65 miles per hour. Jane drove the remainder of the trip and averaged 55 miles per hour. For what length of the time did Dave drive?

Answer by Nate(3500)

Dave and Jane drove a total of 935 miles in 15.2 hours. Dave drove the first part of the trip and averaged 65 miles per hour. Jane drove the remainder of the trip and averaged 55 miles per hour. For what length of the time did Dave drive?
Distance = Rate * Time
~ Dave ~
Rate = 65mph
Time = t
Distance = 65t
~ Jane ~
Rate = 55mph
Time = 15.2 - t
Distance = 836 - 55t
The sum of their distances = The total distance
65t + 836 - 55t = 935
10t = 99
t = 9.9
Dave ~> 9.9 hours
Jane ~> 5.3 hours